OpenStudy (anonymous):

What is the area of the trapezoid. ( my answer is 88in2) other answers: 108in2 and 216in2

OpenStudy (anonymous):

A = 1/2(b1 + bs)h where b1 = 17 b2 = 10 h = 8 Put those in and solve for a :3

OpenStudy (anonymous):

b2**

OpenStudy (anonymous):

like adding them together or timesing , im doing it now on calculator

OpenStudy (anonymous):

you add 17 and 10 and then multiply 1/2 and 8

OpenStudy (mathmale):

Please note that when you add 17 and 10 together and divide the resulting sum by 2, you are finding the average length of the trapezoid. Multiplying that average length by the width (8) gives you the area of the trapezoid.

OpenStudy (anonymous):

well for (1/2) h (b1 + b2) i just found this- >2A/h - b2 = b1

OpenStudy (anonymous):

1/2(17 + 10)8

OpenStudy (anonymous):

so many numbers @.@

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

whats a trap owo

OpenStudy (anonymous):

I tried to simplify but had no luck, but was my answer wrong? @ganeshie8

OpenStudy (anonymous):

what they were saying earlier got me 26, and so I did this, 26 multiplied by 1/2 and 8 and got 21

OpenStudy (anonymous):

@kevkool97

OpenStudy (anonymous):

108 in.

OpenStudy (anonymous):

A=(a+b)h/2 just substitute and see wat u get...

OpenStudy (anonymous):

ah thnks